# Acquiring Better Load Estimates by Combining Anomaly and Change-point Detection in Power Grid Time-series Measurements

### v1

@@ 2.3.1 Base Detectors
Under the assumption that the data is stationary we can use classical statistical process control, or SPC, methods Oakland (2007) in order to detect anomalies and switch events. In traditional SPC one assumes a process which is stationary within a chosen time frame or segment. Then, the user defines certain lower and upper control limits, typically 2 or 3 standard deviations away from the mean of the segment. In our case, the control limits will be optimized as either symmetric or asymmetric thresholds, rather than using the traditional statistical approach. When a time point falls outside of these control limits, it is flagged as out-of-control, in this case meaning anomalous. How SPC is applied to each difference vector $\boldsymbol{\mathbf{\delta}}\in\Delta$ resulting from preprocessing is described in Algorithm 2
@@
One of the most commonly used machine learning methods for anomaly detection is the isolation forest, or IF Liu et al. (2008). An isolation forest on one-dimensional data effectively produces a density estimate by randomly splitting subsets of the data. We consider two distinct ways of applying isolation forests on the data: one where we apply a single forest per station difference vector, see Algorithm 3, and one where we scale and concatenate all difference vectors for training the isolation forest and apply that isolation forest on each difference vector, see Algorithm 4.
@@
In either procedure, we need to set the hyperparameters of the isolation forest. In the case where a single isolation forest is applied, we also consider the quantile boundaries for scaling. Furthermore, the choice between fitting a single forest or a forest per station is treated as a hyperparameter.
@@
Binary segmentation is a change point detection algorithm able to find multiple change points in a given time series Scott and Knott (1974); Bai (1997). We use binary segmentation as a change point algorithm because it is found to be one of the best performing algorithms on real-world univariate time series data Van den Burg and Williams (2020); Schäfer et al. (2021); Ermshaus et al. (2023). It finds change points by recursively partitioning the time series into two parts, forming a binary tree. A split occurs at the optimal break point. This break point is found by first calculating the cost $c_{\textrm{total}}$ of the entire segment using a chosen cost function. Then the costs of the two subsegments, $c_{\textrm{left}}$ and $c_{\textrm{right}}$, are calculated for each possible break point using the same cost function. The optimal break point is then found by selecting the one for which the gain, $g=c_{\textrm{total}}-c_{\textrm{left}}-c_{\textrm{right}}$ is maximized. This procedure is repeated until the gain for a split is below a user-defined threshold called the penalty $p$. In our experiments, we consider two penalties, a linear and a L1 penalty Van den Burg and Williams (2020), the scaling of which depends on a user defined $\beta$ parameter.
We use binary segmentation in order to generate scores by performing the following procedure explained in Algorithm 5
@@
11:for $b_{\textrm{begin}},b_{\textrm{end}}\in\boldsymbol{\mathbf{b}}$ do
12:     $\boldsymbol{\mathbf{t}}\leftarrow(\boldsymbol{\mathbf{\tilde{z}}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}$triangleright Get segment
13:     $(\boldsymbol{\mathbf{z}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}\leftarrow\textrm{mean}(\boldsymbol{\mathbf{t}})-r$triangleright Get difference between segment and reference value
14:end for
This procedure has a large number of optimizable hyperparameters resulting from the binary segmentation algorithm, the strategy for determining the reference point, and with which quantiles scaling should be applied. We have compared 4 different strategies for determining the reference point: 1) comparing to the mean of the entire scaled station difference vector, 2) comparing to the median of the entire scaled station difference vector, 3) comparing to the mean of the longest segment, and 4) comparing to the median of the longest segment.
@@ 2.3.2 Ensembles
The different base detection algorithms have various strengths and weaknesses. Specifically, binary segmentation is good at detecting long events, while SPC and IF are good at detecting shorter events. In order to leverage the strengths of multiple complementary methods, we employ different ensembling techniques to combine base detection algorithms. We specifically compare naive ensembling through combining predictions directly, combining algorithms optimized on detecting events of different lengths, and sequential ensembles, where we apply binary segmentation and apply SPC or IF only on the “normal” segments. A more detailed description of each ensembling strategy can be found below. For all ensembles, we look at a combination of binary segmentation and either IF or SPC.
@@ 2.4.2 Validation
In order to optimize thresholds, hyperparameters and to compare models, we split our original dataset consisting out of $180$ stations into 3 equal parts of 60 stations each, creating a training, validation, and test dataset. This splitting procedure was done in a stratified manner such that the three splits have a more or less equal distribution of event lengths. The stations were divided by aiming to have an equal number of events for each category present in each dataset. As the computational complexity of evaluating all possible combinations is too high, this was done using a greedy approach, where stations were divided starting with the longest, and least frequent, event length category, and ending with the shortest and most frequent event length category. All stations with no events were subsequently divided amongst the datasets in order to create 3 equally large datasets. The number of events, as well as the label “1” counts, can be found in Table 1.
@@
To get an estimate of the reliability of each method, we perform bootstrapping Efron and Tibshirani (1994). We do so by resampling the test set stations with replacement 10000 times. In this way, we acquire both a bootstrapped mean and standard deviation for the precision, recall, and F1.51.51.51.5 metrics.
@@ 2.4.3 Threshold Optimization
All methods we have applied produce scores $\boldsymbol{\mathbf{z}}\in Z$. These can either be positives scores, where a higher value indicates a higher likelihood for a sample to be an anomaly or switch event according to the model, or scores centered around 0, where a greater distance to 0 indicates a higher likelihood. Of the presented methods, isolation forest has a purely positive score, while the statistical process control and binary segmentation scores are centered around 0. These scores are thresholded to yield a label prediction vector $\boldsymbol{\mathbf{\tilde{y}}}\in\tilde{Y}$.
We discern between two strategies for optimizing thresholds: a one-sided or symmetrical approach, $\boldsymbol{\mathbf{\tilde{y}}}=\textrm{thresholdScores}(\textrm{abs}(\boldsymbol{\mathbf{z}}),\theta^{\textrm{symmetrical}})$, and a two-sided or asymmetrical approach, $\boldsymbol{\mathbf{\tilde{y}}}=\textrm{thresholdScores}(\boldsymbol{\mathbf{z}},\theta^{\textrm{lower}},\theta^{\textrm{upper}})$. The symmetrical approach can be used on both positive and zero-centered scores, whereas the asymmetrical approach can only be used on the zero-centered scores.
@@ 2.5 Implementation and Reproducibility
Our analysis has been done in the Python programming language, specifically version 3.10.0. Many of our calculations rely on NumPy Harris et al. (2020) and Pandas The Pandas development team (2023). We furthermore make use of the RupturesTruong et al. (2018) package for binary segmentation, and use the Scikit-learn Pedregosa et al. (2011) package for scaling procedures, as well as applying the isolation forest. Visualizations were made using the Seaborn Waskom (2021) and Matplotlib Hunter (2007) packages. In order to reproduce all our experiments, we have provided access to a public GitHub repository [^1] The Git repository can be found at: https://github.com/RoelBouman/StormPhase2. This reproduces the data splitting procedure, all experiments, including hyperparameter optimization, as well as producing all figures and tables in this paper. All data is provided through Alliander [^2] The data repository can be found at https://www.liander.nl/over-ons/open-data (At time of review, this data is not yet publicly available). An overview of all evaluated, and optimal, hyperparameters can be found in A.
@@ 3 Results
Figure 6: Bar plot of the results of each method in terms of F1.51.51.51.5 per length category. The height of each bar indicates the average F1.51.51.51.5 score over the bootstrap iterations. The error bars indicate the standard deviation resulting from the bootstrap resampling.
@@
We find, as expected, that IF and SPC perform similarly across the board. They are able to detect short anomalies/events better than binary segmentation, achieving a F1.51.51.51.5 score of approximately $0.2$ for both the “15 minutes to 6 hours”, and “6 hours to 3 days” event length categories. Perhaps surprisingly, these methods are able to detect some of the longer switch events as well, achieving just below $0.3$ F1.51.51.51.5 scores. The relatively good performance on these longer events can be easily explained. Many of the longer switch events have a larger distance to the $\boldsymbol{\mathbf{\delta}}$ median than normal data. SPC is therefore able to detect these events based on this distance, while IF identifies them because these switch events occupy a lower density region where they are not masked by normal data. We find that even when ensembled in any of the three tested manners, IF offers no significant performance boost over SPC. Due to the ease of interpretability of SPC, it is preferable over IF.
@@
Figure 7: Bar plot of the results of each method in terms of recall per length category. The height of each bar indicates the average recall score over the bootstrap iterations.
Sequential ensembles seemingly do not suffer from this increase in false positives as much. We can see that sequential ensembles nearly match the performance of the individual SPC and IF detectors in the single categories where they perform best, while the performance in the “3 to 42 days” category is significantly better than any base detector or other ensemble. The performance in the longest category is comparable to that of the other ensembles. Sequential ensembles indeed outperform any base detector or other ensembling method. While intuitively one might think the performance increase in the shorter categories should be similar to other ensembling strategies, it seems that due to the sequential nature, we can optimize more on precision than is the case in other ensembles. This can also be observed in Figure 7, where we see that the recall of the sequential models is lower in the longest category, while being similar in the other three. This means that all performance increases from naive or DOC ensembles to sequential can be attributed to a higher precision.

### v3

@@ 2.3.1 Base Detectors
Under the assumption that the data is stationary we can use classical statistical process control, or SPC, methods (Oakland, 2007) in order to detect anomalies and switch events. In traditional SPC one assumes a process which is stationary within a chosen time frame or segment. Then, the user defines certain lower and upper control limits, typically 2 or 3 standard deviations away from the mean of the segment. In our case, the control limits will be optimized as either symmetric or asymmetric thresholds, rather than using the traditional statistical approach. Additionally, due to the known presence of anomalies in the data, we also look at the distance to the median, rather than to the mean. We will explicitly refer to the optimized control limit as “threshold” from here on out. When a time point falls outside of these control limits, it is flagged as out-of-control, in this case meaning anomalous. How SPC is applied to each difference vector $\boldsymbol{\mathbf{\delta}}\in\Delta$ resulting from preprocessing is described in Algorithm 4
@@
One of the most commonly used machine learning methods for anomaly detection is the isolation forest, or IF (Liu et al., 2008). The isolation forest is known as one of the best state-of-the-art anomaly detectors on real-valued static data (Bouman et al., 2024). An isolation forest on one-dimensional data effectively produces a density estimate by randomly splitting subsets of the data. We consider two distinct ways of applying isolation forests on the data: one where we apply a single forest per station difference vector, see Algorithm 5, and one where we scale and concatenate all difference vectors for training the isolation forest and apply that isolation forest on each difference vector, see Algorithm 6. Note that in these algorithms, we use several high-level functions. “fitPredictIsolationForest” fits an isolation forest with $n_{\textrm{estimators}}$ trees on the input $\boldsymbol{\mathbf{\delta}}$ and returns the anomaly scores $\boldsymbol{\mathbf{\tilde{z}}}$ on the same input. “fitIsolationForest” fits an isolation forest with $n_{\textrm{estimators}}$ trees on the input $\boldsymbol{\mathbf{\delta}}$ and returns the fitted model $\gamma$. “predictIsolationForest” returns the anomaly scores $\boldsymbol{\mathbf{\tilde{z}}}^{i}$ calculated over $\boldsymbol{\mathbf{\delta}}_{\textrm{scaled}}^{i}$ given an already fitted isolation forest $\gamma$. Each of these functions is implemented as part of the scikit-learn library (Pedregosa et al., 2011).
@@
In either procedure, we need to set the hyperparameters of the isolation forest. In the case where a single isolation forest is applied, we also consider the quantile boundaries for scaling. Furthermore, the choice between fitting a single forest or a forest per station is treated as a hyperparameter. We perform rescaling of the scores so that the algorithms deems the most anomalous samples to have the highest scores. These definitions differ between different algorithms in literature, but we chose this definition to be analogous to SPC, allowing for a similar threshold optimization strategy.
@@
Binary segmentation is a change point detection algorithm able to find multiple change points in a given time series (Scott and Knott, 1974; Bai, 1997). We use binary segmentation as a state-of-the-art change point algorithm because it is found to be one of the best performing algorithms on real-world univariate time series data (Van den Burg and Williams, 2020; Schäfer et al., 2021; Ermshaus et al., 2023). It finds change points by recursively partitioning the time series into two parts, forming a binary tree. A split occurs at the optimal break point. This break point is found by first calculating the cost $c_{\textrm{total}}$ of the entire segment using a chosen cost function. Then the costs of the two subsegments, $c_{\textrm{left}}$ and $c_{\textrm{right}}$, are calculated for each possible break point using the same cost function. The optimal break point is then found by selecting the one for which the gain, $g=c_{\textrm{total}}-c_{\textrm{left}}-c_{\textrm{right}}$ is maximized. This procedure is repeated until the gain for a split is below a user-defined threshold called the penalty $p$. In our experiments, we consider two penalties, a linear and a L1 penalty (Van den Burg and Williams, 2020), the scaling of which depends on a user defined $\beta$ parameter.
We use binary segmentation in order to generate scores by performing the following procedure explained in Algorithm 7. The “findReferenceValue” function is described in more detail in Algorithm 8. The “findBreakpointsBinarySegmentation” is a call to the high-level “ruptures” Python library (Truong et al., 2018) which takes the scaled vector $\boldsymbol{\mathbf{\tilde{z}}}$ and segments it according to the well-known binary segmentation algorithm with hyperparameters $\beta,C,l$, and $j$.
@@
11:$b_{\textrm{begin}}=1$triangleright indexing begins at 1
12:for $b_{\textrm{end}}\in\boldsymbol{\mathbf{b}}$ do
13:     $\boldsymbol{\mathbf{t}}\leftarrow(\boldsymbol{\mathbf{\tilde{z}}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}$triangleright Get segment
14:     $(\boldsymbol{\mathbf{z}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}\leftarrow\textrm{mean}(\boldsymbol{\mathbf{t}})-r$triangleright Get difference between segment and reference value
15:     $b_{\textrm{begin}}=b_{\textrm{end}}$
16:end for
Algorithm 8 findReferenceValue
1:Input: Scaled difference vector $\boldsymbol{\mathbf{\tilde{z}}}$, and breakpoints resulting from binary segmentation $\boldsymbol{\mathbf{b}}$
2:Hyperparameters: Reference point strategy reference_point
3:Output: Reference point value $r$
4:
[…]
6:     $r=\textrm{mean}(\boldsymbol{\mathbf{\tilde{z}}})$
[…]
8:     $r=\textrm{median}(\boldsymbol{\mathbf{\tilde{z}}})$
[…]
10:     $s_{\max}=0$
11:     $b_{\textrm{begin}}=1$triangleright indexing begins at 1
12:     for $b_{\textrm{end}}\in\boldsymbol{\mathbf{b}}$ do
13:         $\boldsymbol{\mathbf{t}}\leftarrow(\boldsymbol{\mathbf{\tilde{z}}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}$triangleright Get segment
14:         $s_{i}\leftarrow b_{\textrm{end}}-b_{\textrm{begin}}$triangleright Get segment size
15:         if $s_{i}>s_{\max}$ then
16:              $s_{\max}=s_{i}$
17:              $r=\textrm{mean}(\boldsymbol{\mathbf{t}})$
18:         end if
19:         $b_{\textrm{begin}}=b_{\textrm{end}}$
20:     end for
[…]
22:     $s_{\max}=0$
23:     $b_{\textrm{begin}}=1$triangleright indexing begins at 1
24:     for $b_{\textrm{end}}\in\boldsymbol{\mathbf{b}}$ do
25:         $\boldsymbol{\mathbf{t}}\leftarrow(\boldsymbol{\mathbf{\tilde{z}}}_{i})_{b_{\textrm{begin}}\leq i\leq b_{\textrm{end}}}$triangleright Get segment
26:         $s_{i}\leftarrow b_{\textrm{end}}-b_{\textrm{begin}}$triangleright Get segment size
27:         if $s_{i}>s_{\max}$ then
28:              $s_{\max}=s_{i}$
29:              $r=\textrm{median}(\boldsymbol{\mathbf{t}})$
30:         end if
31:         $b_{\textrm{begin}}=b_{\textrm{end}}$
32:     end for
33:end if
This procedure has a large number of optimizable hyperparameters resulting from the binary segmentation algorithm, the strategy for determining the reference point, and with which quantiles scaling should be applied. We have compared 4 different strategies for determining the reference point: 1) comparing to the mean of the entire scaled station difference vector (“mean”), 2) comparing to the median of the entire scaled station difference vector(“median”), 3) comparing to the mean of the longest segment (“longest_mean”), and 4) comparing to the median of the longest segment (“longest_median”).
@@ 2.3.2 Ensembles
The different base detection algorithms have various strengths and weaknesses. Specifically, binary segmentation is good at detecting long events, while SPC and IF are good at detecting shorter events. Binary segmentation more easily detects long events as it can compare the distance of an entire segment to the distance of a different segment, thus being more sensitive to small changes over a long time period. SPC and IF consider each timepoint individually, thus only detecting large changes without considering the time component. In order to leverage the strengths of multiple complementary methods, we employ different ensembling techniques to combine base detection algorithms. We specifically compare naive ensembling through combining predictions directly, combining algorithms optimized on detecting events of different lengths, and sequential ensembles, where we apply binary segmentation and apply SPC or IF only on the “normal” segments. A more detailed description of each ensembling strategy can be found below. For all ensembles, we look at a combination of a change point detector, binary segmentation, and an anomaly detector, which is either SPC or IF. As both SPC and IF are designed to find singular anomalies, we see them as complementary to binary segmentation, and will specifically compare SPC to IF.
@@ 2.4.2 Validation
In order to optimize thresholds, hyperparameters and to compare models, we split our original dataset consisting out of $180$ stations into 3 equal parts of 60 stations each, creating a training, validation, and test dataset. This splitting procedure was done in a stratified manner such that the three splits have a more or less equal distribution of event lengths. The stations were divided by aiming to have an equal number of events for each category present in each dataset. The computational complexity of evaluating all possible combinations, like is done in available cross-validation software, is too high. We have therefore used a greedy approach to produce a stratified split of the data, where stations were divided starting with the longest, and least frequent, event length category, and ending with the shortest and most frequent event length category. All stations with no events were subsequently divided amongst the datasets in order to create 3 equally large datasets. The number of events, as well as the label “1” counts, can be found in Table 1.
@@
To get an estimate of the reliability of each method, we perform bootstrapping (Efron and Tibshirani, 1994). We do so by resampling the test set stations with replacement 10,000 times. In this way, we acquire both a bootstrapped mean and standard deviation for the precision, recall, and F1.51.51.51.5 metrics.
@@ 2.4.3 Threshold Optimization
All methods we have applied produce scores $\boldsymbol{\mathbf{z}}\in Z$. These can either be positives scores, where a higher value indicates a higher likelihood for a sample to be an anomaly or switch event according to the model, or scores centered around 0, where a greater distance to 0 indicates a higher likelihood. Of the presented methods, isolation forest has a purely positive score, while the statistical process control and binary segmentation scores are centered around 0. These scores are thresholded to yield a label prediction vector $\boldsymbol{\mathbf{\tilde{y}}}\in\tilde{Y}$. This is done by applying a one-sided or symmetrical approach, $\boldsymbol{\mathbf{\tilde{y}}}=\textrm{thresholdScores}(\textrm{abs}(\boldsymbol{\mathbf{z}}),\theta^{\textrm{symmetrical}})$, or a two-sided or asymmetrical approach, $\boldsymbol{\mathbf{\tilde{y}}}=\textrm{thresholdScores}(\boldsymbol{\mathbf{z}},\theta^{\textrm{lower}},\theta^{\textrm{upper}})$. The symmetrical approach can be used on both positive and zero-centered scores, whereas the asymmetrical approach can only be used on the zero-centered scores. Pseudocode for the one-sided approach can be found in Algorithm 9, and for the two-sided approach in Algorithm 10.
Algorithm 9 thresholdScores (one-sided)
1:Input: A score vector $\boldsymbol{\mathbf{z}}$, where a higher $z_{i}$ indicates a higher likelihood of being an anomaly/switch event
2:Hyperparameters: A threshold $\theta^{\textrm{symmetrical}}$
3:Output: Predicted label vector $\boldsymbol{\mathbf{\tilde{y}}}$
4:for $z_{i}\in\boldsymbol{\mathbf{z}}$ do
5:
$\tilde{y}_{i}\leftarrow\begin{cases}\textrm{1},&\text{if }z_{i}\geq\theta^{\textrm{symmetrical}}\\ \textrm{0},&\text{otherwise}\end{cases}$
6:end for
Algorithm 10 thresholdScores (two-sided)
1:Input: A score vector $\boldsymbol{\mathbf{z}}$, where a high or low $z_{i}$ indicates a higher likelihood of being an anomaly/switch event
2:Hyperparameters: A lower threshold $\theta^{\textrm{lower}}$, and upper threshold $\theta^{\textrm{upper}}$
3:Output: Predicted label vector $\boldsymbol{\mathbf{\tilde{y}}}$
4:for $z_{i}\in\boldsymbol{\mathbf{z}}$ do
5:
$\tilde{y}_{i}\leftarrow\begin{cases}\textrm{1},&\text{if }z_{i}\geq\theta^{\textrm{upper}}\\ \textrm{1},&\text{else if }z_{i}<\theta^{\textrm{lower}}\\ \textrm{0},&\text{otherwise}\end{cases}$
6:end for
@@ 2.5 Implementation and Reproducibility
Our analysis has been done in the Python programming language, specifically version 3.10.0. Many of our calculations rely on NumPy (Harris et al., 2020) and Pandas (The Pandas development team, 2023). We furthermore make use of the Ruptures(Truong et al., 2018) package for binary segmentation, and use the Scikit-learn (Pedregosa et al., 2011) package for scaling procedures, as well as applying the isolation forest. Visualizations were made using the Seaborn (Waskom, 2021) and Matplotlib (Hunter, 2007) packages. In order to reproduce all our experiments, we have provided access to a public GitHub repository [^1] The Git repository can be found at: https://github.com/RoelBouman/StormPhase2. This reproduces the data splitting procedure, all experiments, including hyperparameter optimization, as well as producing all figures and tables in this paper. All data is provided through Alliander [^2] The data repository can be found at https://www.liander.nl/over-ons/open-data (At time of review, this data is not yet publicly available). An overview of all evaluated, and optimal, hyperparameters can be found in A.
@@ 3 Results
Figure 6: From top to bottom: bar plots of the results of each method per length category for the F1.5, recall, and precision respectively. The height of each bar indicates the average score over the bootstrap iterations. The error bars indicate the standard deviation resulting from the bootstrap resampling.
@@
We find that IF and SPC perform similarly across the board. This is partially to be expected, as they are designed to find short anomalies/events that are rare compared to normal data. IF, however is able to more easily model multimodal distributions and detect anomalies in the presence of multimodality. Since IF does not noticably outperform SPC, it seems that it can not leverage this advantage. This is further corroborated by visualization of typical scaled data, which is mostly unimodal. both IF and SPC are able to detect short anomalies/events better than binary segmentation, achieving a F1.51.51.51.5 score of approximately $0.2$ for both the “15 minutes to 6 hours”, and “6 hours to 3 days” event length categories. Perhaps surprisingly, these methods are able to detect some of the longer switch events as well, achieving just below $0.3$ F1.51.51.51.5 scores. The relatively good performance on these longer events can be easily explained. Many of the longer switch events have a larger distance to the $\boldsymbol{\mathbf{\delta}}$ median than normal data. SPC is therefore able to detect these events based on this distance, while IF identifies them because these switch events occupy a lower density region where they are not masked by normal data. We find that even when ensembled in any of the three tested manners, IF offers no significant performance boost over SPC. Due to the ease of interpretability of SPC, it is preferable over IF.
@@
Sequential ensembles seemingly do not suffer from this increase in false positives as much. We can see that sequential ensembles nearly match the performance of the individual SPC and IF detectors in the single categories where they perform best, while the performance in the “3 to 42 days” category is significantly better than any base detector or other ensemble. The performance in the longest category is comparable to that of the other ensembles. Sequential ensembles indeed outperform any base detector or other ensembling method. While intuitively one might think the performance increase in the shorter categories should be similar to other ensembling strategies, it seems that due to the sequential nature, we can optimize more on precision than is the case in other ensembles. This can also be observed in Figure 6, where we see that the recall of the sequential models is lower in the longest category, while being similar in the other three. This means that all performance increases from naive or DOC ensembles to sequential can be attributed to a higher precision, which is confirmed by Figure 6.
@@
Due to the high performance of our ensemble method for load estimation it has been adopted for use within Alliander. Previously, all filtering and load estimates were done by hand. This was a time-consuming process, taking a month of time for several full-time employees. This was done once a year to acquire the load estimates for future planning and operations management. Earlier variations of the presented methodology have been in use since 2021, fully replacing the manual, time-consuming, process. In addition to replacing time-consuming manual labor of domain experts whose expertise is better utilized elsewhere, our methodology allows for easier updating of load estimates, which can now be generated on-the-fly. We hope that by open sourcing our methodology, code, and data other distribution system operators can benefit similarly.